OpenStudy (bethgrace96):

please answer for a medal! Record the next three terms for each sequence. then write an explicit function for each sequence. u0=45, un=1.2un-1 where n is greater than or equal to 1

Nnesha (nnesha):

un and u 0 ??

Nnesha (nnesha):

\[\large\rm u_o = 45 ~~~~u_n =1.2_{n-1}-1\] looks like this ?

OpenStudy (bethgrace96):

yes. but the second one has a U after the 1.2

Nnesha (nnesha):

ohh i forgot sorry \[\large\rm u_o = 45 ~~~~u_n =1.2u_{n-1}-1\] so is it geometric or arithmetic ??

OpenStudy (bethgrace96):

I don't know, thats all the problem said

Nnesha (nnesha):

you can find out by looking at the \[\large\rm u_n =1.2u_{n-1}\] do you know the formulas for geometric and arithmetic sequence ?

OpenStudy (bethgrace96):

Its asking for the next three terms. please just give answers, and re-read my question

Nnesha (nnesha):

i know what they are asking for tell me what is geometric sequence what is arithmetic sequence ? do you know the formulas if you don't know what r or d or if you don't know the formula for geometric or arithmetic seq you can't find 3 terms

OpenStudy (bethgrace96):

http://www.mathwords.com/a/arithmetic_sequence.htm

Nnesha (nnesha):

nice link..! recursive formula for geometric sequence \[\huge\rm a_n = a_{n-1} \times r \] and recursive formula for arithmetic seq is \[\huge\rm a_n = a_{n-1} + d\]

OpenStudy (bethgrace96):

awesome so whats the answer

Nnesha (nnesha):

Geometric sequence: the ratios of the consecutive terms are the same ; geometric when you multiply previous term to get the next term Arithmetic sequence: when the difference between consecutive terms are the same when you add or subtract previous term to get the next one

OpenStudy (bethgrace96):

i don't have much time whats the answer

Nnesha (nnesha):

\[\large\rm u_n =1.2u_{n-1}\] now after reading that info what you think is it geometric or arithmetic ??

OpenStudy (bethgrace96):

arithmetic

Nnesha (nnesha):

how ??

OpenStudy (bethgrace96):

oh wait no geometric now whats the next step

Nnesha (nnesha):

why is that geometric ??

OpenStudy (bethgrace96):

because multi

OpenStudy (bethgrace96):

please just work it out so i can see it!!!!!!!!!!

Nnesha (nnesha):

Geometric = ratio = multiplication Arithmetic = difference = + sign

OpenStudy (bethgrace96):

okay great can u answer the question??? this is taking forever.

Nnesha (nnesha):

yep that's right geometric \[\large\rm u_n =1.2u_{n-1}\] so r = 1.2 explicit formula for geometric sequence is \[\large a_n = a_1 ( r)^{n-1}\] r= ratio n = term that you have to find (1 ,2 ,3) replace variables by their values

Nnesha (nnesha):

is that \[\large\rm u_o ~~~or ~~ u_1 ???\]

OpenStudy (bethgrace96):

0

OpenStudy (bethgrace96):

so whats the answer

OpenStudy (bethgrace96):

helloo????????????

Nnesha (nnesha):

 to find first term replace u_o by 45 r by 1.2 and n =1 and solve for u_1